A ball is dropped at t=3s. FindA) Instantaneous speed at t=18sB) avg. Speed at t=20s

Grade:11

1 Answers

sonika p
6 years ago
since, ball is dropped at t = 3s, initial speed = 0
v= gt,
g=9.8m/s2
  1. instantaneous speed – = 9.8*18 = 176.4m/s
  2. average speed = 0.5*9.8*(20+3) = 112.7m/s